OpenStudy (anonymous):

Which of the following functions will produce a graph that is a horizontally compressed version of the graph of y = x2?

OpenStudy (kira_yamato):

To have a compressed graph, say scale factor = n, \[n \in \mathbb{R} , n>1\] Replace x by nx y = (nx)^2

OpenStudy (kira_yamato):

I mean scale factor = 1/n. Sorry...